archive-es.com » ES » O » OVH.ES

Total: 788

Choose link from "Titles, links and description words view":

Or switch to "Titles and links view".
  • OVH Tasks  
    FS 6164 MTU change on the collect Closed 2011 12 15 2195 ADSL SDSL Modernization Backend Core FS 6166 Eligibility per address Closed 2011 12 15 2219 ADSL SDSL Modernization Backend Core FS 6188 LNS update Closed 2011 12 30 2285 ADSL SDSL Modernization Backend Core FS 6249 Lns IPV6 Closed 2012 01 13 2312 ADSL SDSL Modernization Backend Core FS 6280 Support of the DNSSec validation on the DNS cache Closed 2012 01 18 2525 ADSL SDSL Modernization Backend Core FS 6489 Marseille collection system Closed 2012 03 14 2654 ADSL SDSL Modernization Backend Core FS 6616 Lille dslam loop Closed 2012 04 20 2656 ADSL SDSL Modernization Backend Core FS 6614 Marseille optical loop set up of CAV13 Closed 2012 04 19 2903 ADSL SDSL Modernization Backend Core FS 6864 Marseille securisation Closed 2012 06 11 3046 ADSL SDSL Modernization Backend Core FS 6999 Marseille CAV13 Closed 2012 07 12 3071 ADSL SDSL Modernization Backend Core FS 7024 LNS Bordeaux update Closed 2012 07 17 3084 ADSL SDSL Modernization Backend Core FS 7037 Paris STC92 CVL92 ROB92 VER78 VIR78 Closed 2012 07 18 3234 ADSL SDSL Modernization Backend Core FS 7185 LNS updating for Paris Roubaix Lyon and Marseille Closed 2012 08 17 3429 ADSL SDSL Modernization Backend Core FS 7374 Bordeaux AUG33 MQ233 BT733 Closed 2012 09 27 3543 ADSL SDSL Modernization Backend Core FS 7485 switches Paris Closed 2012 10 17 3590 ADSL SDSL Modernization Backend Core FS 7533 Upgrade switchs Closed 2012 10 30 3709 ADSL SDSL Modernization Backend Core FS 7651 lil 3 4 n5 Closed 2012 11 20 3727 ADSL SDSL Modernization Backend Core FS 7669 Update Switchs Closed 2012 11 22 3825 ADSL SDSL Modernization Backend Core FS 7767 Improving data Server Closed 2012 12 13 3873 ADSL SDSL Modernization Backend Core

    Original URL path: http://status.ovh.es/?order=type&project=0&tasks=&string=&type=&sev=&dev=&due=&cat=&status=&order2=proj&sort=asc&sort2=asc&date=&project=&sort=desc&order2=proj&sort2=asc (2015-10-15)
    Open archived version from archive


  • OVH Tasks  
    Incident Category Backend Core Status CLOSED Percent Complete Details We are facing right now a problem of the LNS of Roubaix We are suspending all the LNS change requests to Roubaix The operations will be re runned later when the Roubaix LNS will be stable We thank you for your understanding Date Friday 08 April 2011 15 40PM Reason for closing Done Additional comments about closing The LNS of roubaix

    Original URL path: http://status.ovh.es/?do=details&id=1309 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Percent Complete Details We found a problem on the SDSL card of the DSLAM We must intervene on the equipment to proceed for the replacement of the card We will begin the intervention at 9 pm A few interrruptions are

    Original URL path: http://status.ovh.es/?do=details&id=1407 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    C A technician is underway to replace the equipment An interuption of a few minutes is expected during maintenance Date Tuesday 14 June 2011 19 34PM Reason for closing Done Comments 3 Comment by OVH Tuesday 14 June 2011 19 32PM The first optical bloc faces itself problems A thechnician is on the road to replace it Comment by OVH Tuesday 14 June 2011 19 32PM We are replacing the

    Original URL path: http://status.ovh.es/?do=details&id=1533 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Category Backend Core Status CLOSED Percent Complete Details There is a problem on the NRA of Lesquin Date Tuesday 21 June 2011 03 30AM Reason for closing Done Additional comments about closing The optic is UP The network is operational Comments 2 Comment by OVH Tuesday 21 June 2011 01 16AM A team is being informed to deal with NRA problem Comment by OVH Tuesday 21 June 2011 01 18AM

    Original URL path: http://status.ovh.es/?do=details&id=1552 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Complete Details There is an electrical problem in the region of Ain and Rhone Around twenty DSLAM are disrupted A generator is supplying the sites Date Tuesday 23 August 2011 10 49AM Reason for closing Done Additional comments about closing

    Original URL path: http://status.ovh.es/?do=details&id=1731 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    is always passing we will verify will it Date Tuesday 30 August 2011 13 58PM Reason for closing Done Comments 2 Comment by OVH Tuesday 30 August 2011 13 55PM Part of the routing goes in soft on our equipment in LIB59 which also impacts the other NRAs of this loop SEC59 FAC59 BOI59 BLE59 We need to reconfigure the equipment which will require a reload DSL traffic on these

    Original URL path: http://status.ovh.es/?do=details&id=1761 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Complete Details SFR is performing works on the fiber connecting the NRA from WAS59 to SAN59 Customers located on the NRA WAS59 are temporarily cut of the network The maintenance is scheduled until 6 00 Date Tuesday 13 September 2011

    Original URL path: http://status.ovh.es/?do=details&id=1804 (2015-10-15)
    Open archived version from archive



  •